At 25 °C, only 0.0870 mol of the generic salt AB is soluble in 1.00 L of water. What is the K sp of the salt at 25 °C? AB ( s ) − ⇀ ↽ − A + ( aq ) + B − ( aq ) K sp =
Answer #1

Molar solubility(s) = 0.0870/1.00 = 0.0870 M

Ksp = s​​​​​​2 = 7.57×10-3

So, Ksp is 7.57×10-3